OpenStudy (anonymous):

Let ρAl represent the density of aluminum and ρFe that of iron. Find the radius of a solid aluminum sphere that balances a solid iron sphere of radius rFe on an equal-arm balance. (Use any variable or symbol stated above as necessary.)

OpenStudy (anonymous):

in the form of rAl= ( answer basically has to be a formula over a formula)

OpenStudy (anonymous):

\[\frac{4}{3} * \pi (rAL)^{3} *g= \frac{4}{3} * \pi (rFe)^{3} *g\]

OpenStudy (anonymous):

your balancing the m1*g=m2*g

OpenStudy (anonymous):

woops forgot the density, its volume * density *gravity .....on both sides

OpenStudy (anonymous):

\[ rAL=\sqrt[3]{rFE^{3} * \frac{\rho FE}{\rho Al}}\]

OpenStudy (anonymous):

Right on Point, Thank You

OpenStudy (anonymous):

you can simplify and write rFE on the outside of the cube root